# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 3.1 Integrating out heavy flavors
In this paper, we integrate out heavy quarks and leptons at, or close to, their mass threshold. In particular, we integrate out the $b$ quark at $\mu_{b}=m^{\overline{\rm MS}}_{b}=4.18$ GeV, and the charm quark at $\mu_{c}=2$ GeV. We thus do not consider nonperturbative matrix elements of operators with $b$ and $c$ quarks between nucleons or nucleons and pions, but take their effects into account in perturbation theory, via renormalization group mixing or matching at the threshold. For the $b$ quark, with $\alpha_{s}(m^{\overline{\rm MS}}_{b})$ well in the perturbative regime, this approach is fully justified. Since the $c$ quark lies at the boundary between the perturbative and nonperturbative regimes, in this case one could pursue the alternative approach of not integrating out the $c$, and evaluating charm matrix elements in Lattice QCD. While most Lattice QCD collaborations use actions with dynamical charm quarks, calculations of nucleon charm matrix elements are very preliminary in the case of quark bilinears Alexandrou:2019brg , and non-existing in the case of more complicated operators, such as the quark chromoelectric dipole moment or four-fermion operators. One would thus have to rely on models, which come with uncontrolled uncertainties. As Lattice QCD calculations improve, it will become interesting to compare the perturbative and nonperturbative treatment of heavy quarks.
@@ 3.2 Leptonic EDMs
For the electron, muon and $\tau$ EDMs, there is a direct connection between the WET dipole operators $C_{e\gamma}$ and the electric dipole moment. In addition, the semileptonic tensor operators $\left[C^{T,RR}_{eq}\right]$ with light quark flavors induce charged lepton dipole operator non-perturbatively Dekens:2018pbu ; Aebischer:2021uvt , with a contribution proportional to the correlator of the vector and tensor currents $\Pi_{VT}({\bf q})$ Cirigliano:2021img
@@
$=-1.3\cdot 10^{-16}\left[C_{e\gamma}-2.5\cdot 10^{-5}\,{\rm{TeV}}\left(2\left[C^{T,RR}_{eu}\right]_{1111}-\left[C^{T,RR}_{ed}\right]_{1111}-\left[C^{T,RR}_{ed}\right]_{1122}\right)\right]_{\ell\ell}\,{\rm{TeV}}\,e\,{\rm cm}.$ (21)
In Eq. (3.2), and in the rest of this section, the WET couplings are understood to be evaluated at the scale of $\mu=2$ GeV. Here we use a resonance model for $\Pi_{VT}$ and write Knecht:2001xc ; Mateu:2007tr ; Cata:2008zc ; Cirigliano:2021img
@@
with the quark condensate $B_{0}=2.7$ GeV at the renormalization scale $\mu=2$ GeV, $F_{\pi}=92.2$ MeV and $M_{\rho}=0.770$ GeV. The above estimate is based on large $N_{C}$ considerations, and on the truncation of the resonance spectrum to the first state, and it is therefore affected by large theoretical uncertainties, which was estimated to be at least 50% Knecht:2001xc ; Mateu:2007tr ; Cata:2008zc ; Cirigliano:2021img .
@@ 3.3 ThO, YbF, and HfF Frequencies
The best limit on the electron EDM currently comes from experiments with polar molecules. In this case, the observables are $P$ and $T$-odd frequency shifts $\omega$, which receive contributions from the electron EDM and semileptonic scalar operators. The frequency shifts are expressed in terms of an isospin invariant and an isospin-breaking electron-nucleon coupling, captured by the Lagrangian
$\mathcal{L}=-\frac{G_{F}}{\sqrt{2}}\bar{e}i\gamma_{5}e\,\bar{N}\left(C_{S}^{(0)}+\tau_{3}C^{(1)}_{S}\right)N.$ (24)
The nucleon-level scalar couplings can be expressed in terms of WET four-fermion operators as
@@
where the scaler charge is evaluated at the renormalization scale $\mu=2$ GeV. With these definitions, we can express the frequency measured in the HfF, ThO, and YbF experiments as Degenkolb:2024eve ; Chupp:2017rkp ; Dekens:2018bci
@@ 3.4 Nucleon EDMs
Lattice QCD calculations of the nucleon EDM induced by the quark chromo-electric dipole moment (qCEDM) operators $\left[C_{ug}\right]_{11}$, $\left[C_{dg}\right]_{11,22}$ are at the moment only preliminary Bhattacharya:2023qwf ; Abramczyk:2017oxr ; Kim:2021qae . Currently, the best estimates comes from QCD sum rules Pospelov:2000bw ; Pospelov:2005pr ; Lebedev:2004va ; Fuyuto:2012yf . In Refs. Pospelov:2000bw ; Pospelov:2005pr ; Lebedev:2004va , the $u$ and $d$ qCEDM contribution to the neutron and proton EDMs was estimated to be
@@
with roughly 50% uncertainty. Here $g_{s}$ is the strong coupling constant. Ref. Fuyuto:2012yf finds a larger value and a substantial contribution from the strange qCEDM
@@
Estimates based on the dominance of long-distance contributions mediated by pion-nucleon couplings end up in a similar range Seng:2014pba ; deVries:2016jox , though, of course, they are affected by uncontrolled systematic errors. The above matrix elements are given in the assumption that a PQ mechanism cancels the $\bar{\theta}$ term, including the shift in the $\bar{\theta}$ term induced by the chromoelectric operators Pospelov:2005pr . The scale dependence of $g_{s}$ and of the WET coefficients should in principle be compensated by the scale dependence of the hadronic matrix elements. In the sum rule calculations, such scale dependence is not explicitly tracked, and thus, at the moment, the choice of $\mu$ introduces some arbitrariness. Since the conversion between Lattice QCD schemes and $\overline{\rm MS}$ is known Bhattacharya:2015rsa ; Mereghetti:2021nkt , this issue will be overcome once Lattice QCD calculation will become available. The contribution of the Weinberg three-gluon operator $C_{\tilde{G}}$ has been estimated in Refs. Demir:2002gg ; Haisch:2019bml . The two evaluations are in good agreement, and here we use the results of Ref. Haisch:2019bml
@@
The situation for four-fermion operators is even less well-developed. The matching between gradient flow and $\overline{\rm MS}$ renormalization at one loop has been derived in Ref. Buhler:2023gsg . Robust calculations of the nucleon EDM currently do not exist. Here we will use the expression derived at N2LO in two-flavor chiral perturbation theory Seng:2014pba
@@
with $g_{A}=1.27$. $\kappa_{1,0}$ are the nucleon isovector and isoscalar anomalous magnetic moments, $\kappa_{1}=3.7$, $\kappa_{0}=-0.12$. $\bar{g}_{0}$, $\bar{g}_{1}$, $\bar{d}_{n}$, and $\bar{d}_{p}$ are parity ($P$) and time-reversal ($T$) odd low-energy couplings in the nucleon EFT, which implicitly depend on WET Wilson coefficients. In the following, we will provide a recipe to estimate $\bar{g}_{0,1}$ in the case of four-fermion operators.
@@ 3.4.1 Pion nucleon couplings
$P$ and $T$-odd pion nucleon couplings play an important role for both nucleon and nuclear EDMs. In the case of the nucleon EDM, pion-nucleon couplings induce logarithmically-enhanced contributions Crewther:1979pi , as can be seen in Eqs. (42) and (43). Since they are accompanied by unknown low-energy constants, pion-nucleon couplings cannot completely determine the leading contributions to the nucleon EDM, and thus expressions as Eq. (42) will be eventually superseded by Lattice QCD calculations. Knowing the strength of $\bar{g}_{0,1}$, however, can provide guidance for the chiral and momentum extrapolation Dragos:2019oxn and for the removal of excited state contamination Bhattacharya:2021lol in Lattice QCD calculations. Depending on the chiral properties of the WET operators, pion-nucleon couplings furthermore provide one of leading contribution to the $PT$ odd potential deVries:2012ab ; Maekawa:2011vs ; deVries:2020iea , which then feeds into the calculation of Schiff moments and thus atomic EDMs Haxton:1983dq ; Engel:1999np .
We focus here on the most important $\pi$-N couplings and refer for a more detailed discussion to Ref. deVries:2020iea . We define the $PT$-odd $\pi$-N Lagrangian as
$\mathcal{L}_{\pi NN}=\bar{N}\left[\frac{\bar{g}_{0}}{F_{\pi}}\bm{\tau}\cdot\bm{\pi}+\frac{\bar{g}_{1}}{F_{\pi}}\pi_{0}+\frac{\bar{g}_{2}}{F_{\pi}}\left(\pi_{0}\tau_{3}-\frac{1}{3}\bm{\tau}\cdot\bm{\pi}\right)\right]N,$ (44)
where $\bm{\tau}$ are Pauli matrices in isospin space. The couplings in Eq. (44) break chiral symmetry and are thus mostly important for chiral-symmetry-breaking operators. These usually come in chiral multiplets, with the CP-even components inducing corrections to the baryon spectrum while the CP-odd components contribute to pion-nucleon couplings. At leading order in chiral perturbation theory, chiral symmetry then provides a relation between corrections to the baryon spectrum and $\pi$-N couplings deVries:2012ab , which survives higher order corrections deVries:2015una ; deVries:2016jox . It can be shown that $\bar{g}_{2}$ only arises at higher order in isospin breaking deVries:2012ab . The couplings $\bar{g}_{0}$ and $\bar{g}_{1}$ receive a “direct” contribution, proportional to the shift in the baryon masses induced by dim-6 operators, and “indirect” contributions, arising from “vacuum alignment” Baluni:1978rf , i.e. from imposing the condition that $\pi_{0}$ and $\eta$ mesons do not disappear into the vacuum in the presence of $PT$-odd interactions. We will thus write
@@
For the qCEDM, the pion-nucleon couplings were extracted using QCD sum rules Pospelov:2001ys ,
@@
Again, we are neglected contributions proportional to $\bar{\theta}_{\rm ind}$, which are cancelled if a PQ mechanism is active. In chiral perturbation theory, the indirect part of the coupling is given by deVries:2016jox
@@
The ratio between $\bar{g}_{1}/\bar{g}_{0}$ is determined by the ratio of isoscalar and isovector charges, which, for the value of the sigma term we adopt, is about $17$. The assumption of the dominance of the indirect piece leads to larger values compared to the QCD sum rule calculations, especially in the case $\bar{g}_{1}$. For the qCEDM operators we will just use the sum rule results, while for the four-fermion operators, where only the indirect pieces can be estimated, we will capture these effects by assigning a large uncertainty. Plugging Eqs. (46) and (47) or Eqs. (48) and (49) in the chiral perturbation theory expression for the nucleon EDM we recover estimates that agree in magnitude with Eqs. (37) and (39).
We discussed in some detail various estimates for the qCEDM operators to justify our choices for four-fermion operators, for which complete evaluations of the couplings, either on the Lattice or from QCD sum rules are not available. The structure of the couplings is very similar to qCEDM, with a direct piece that could be extracted from dim-6 corrections to the baryon spectrum and an indirect piece determined by the ratio of vacuum matrix elements Cirigliano:2016yhc ; Dekens:2022gha . The operators constructed in Section 2.1.3 belong to the ${\mathbf{8}\times\mathbf{8}}$, $\mathbf{6}\times\mathbf{\bar{6}}$ and $\mathbf{3}\times\mathbf{\bar{3}}$ representations of the $SU(3)$ flavor group Dekens:2022gha . The indirect components of the pion-nucleon couplings are given by
@@
$C^{(1,0)}_{\mathbf{8}\times\mathbf{8}}$ $=\frac{1}{2}\left({\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1221}+{\rm Im}\left[C^{V1,\,LR}_{dd}\right]_{1221}\right)$ (53)
$C^{(8,0)}_{\mathbf{8}\times\mathbf{8}}$ $={\frac{1}{2}}\left({\rm Im}\left[C^{V8,\,LR}_{uddu}\right]_{1221}+{\rm Im}\left[C^{V8,\,LR}_{dd}\right]_{1221}\right)$ (54)
$C^{(1,1)}_{\mathbf{8}\times\mathbf{8}}$ $={\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1111}+\frac{1}{2}{\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1221}-\frac{1}{2}{\rm Im}\left[C^{V1,\,LR}_{dd}\right]_{1221}$ (55)

### v2

@@ 3.1 Integrating out heavy flavors
In this paper, we integrate out heavy quarks and leptons at, or close to, their mass threshold. In particular, we integrate out the $b$ quark at $\mu_{b}=m^{\overline{\rm MS}}_{b}=4.18$ GeV, and the charm quark at $\mu_{c}=2$ GeV. We thus do not consider nonperturbative matrix elements of operators with $b$ and $c$ quarks between nucleons or nucleons and pions, but take their effects into account in perturbation theory, via renormalization group mixing or matching at the threshold. For the $b$ quark, with $\alpha_{s}(m^{\overline{\rm MS}}_{b})$ well in the perturbative regime, this approach is fully justified. Since the $c$ quark lies at the boundary between the perturbative and nonperturbative regimes, in this case one could pursue the alternative approach of not integrating out the $c$, and evaluating charm matrix elements in Lattice QCD. While most Lattice QCD collaborations use actions with dynamical charm quarks, calculations of nucleon charm matrix elements are very preliminary in the case of quark bilinears Alexandrou:2019brg , and non-existing for more complicated operators, such as the charm quark chromoelectric dipole moment or four-fermion operators. One would thus have to rely on models, which come with uncontrolled uncertainties. As Lattice QCD calculations improve, it will become interesting to compare the perturbative and nonperturbative treatment of heavy quarks.
@@ 3.2 Leptonic EDMs
For the electron, muon, and $\tau$ EDMs, there is a direct connection between the WET dipole operators $C_{e\gamma}$ and the electric dipole moment. In addition, the semileptonic tensor operators $\left[C^{T,RR}_{eq}\right]$ with light quark flavors induce charged lepton dipole operator non-perturbatively Dekens:2018pbu ; Aebischer:2021uvt , with a contribution proportional to the correlator of the vector and tensor currents $\Pi_{VT}({\bf q})$ Cirigliano:2021img
@@
$=-1.3\cdot 10^{-16}\left[C_{e\gamma}-2.5\cdot 10^{-5}\,{\rm{TeV}}\left(2\left[C^{T,RR}_{eu}\right]_{1111}-\left[C^{T,RR}_{ed}\right]_{1111}-\left[C^{T,RR}_{ed}\right]_{1122}\right)\right]_{\ell\ell}\,{\rm{TeV}}\,e\,{\rm cm},$ (21)
with $q_{u}=2/3$ and $q_{d}=q_{s}=-1/3$. In Eq. (3.2), and in the rest of this section, the WET couplings are understood to be evaluated at the scale of $\mu=2$ GeV. Here we use a resonance model for $\Pi_{VT}$ and write Knecht:2001xc ; Mateu:2007tr ; Cata:2008zc ; Cirigliano:2021img
@@
with the quark condensate $B_{0}=2.7$ GeV at the renormalization scale $\mu=2$ GeV, $F_{\pi}=92.2$ MeV and $M_{\rho}=0.770$ GeV. The above estimate is based on large $N_{C}$ considerations, and on the truncation of the resonance spectrum to the first state, and it is therefore affected by large theoretical uncertainties, which were estimated to be at least 50% Knecht:2001xc ; Mateu:2007tr ; Cata:2008zc ; Cirigliano:2021img .
@@ 3.3 ThO, YbF, and HfF Frequencies
The best limit on the electron EDM currently comes from experiments with polar molecules. In this case, the observables are the parity ($P$) and time-reversal ($T$) odd frequency shifts $\omega$, which receive contributions from the electron EDM and semileptonic scalar operators. The frequency shifts are expressed in terms of an isospin invariant and an isospin-breaking electron-nucleon coupling, captured by the Lagrangian
$\mathcal{L}=-\frac{G_{F}}{\sqrt{2}}\bar{e}i\gamma_{5}e\,\bar{N}\left(C_{S}^{(0)}+\tau_{3}C^{(1)}_{S}\right)N,$ (24)
with $N=(p,n)$ the nucleon doublet and $\tau_{3}$ a Pauli matrix in isospin space. The nucleon-level scalar couplings can be expressed in terms of WET four-fermion operators as
@@
where the scalar charge is evaluated at the renormalization scale $\mu=2$ GeV. With these definitions, we can express the frequency measured in the HfF, ThO, and YbF experiments as Degenkolb:2024eve ; Chupp:2017rkp ; Dekens:2018bci
@@ 3.4 Nucleon EDMs
Lattice QCD calculations of the nucleon EDM induced by the quark chromoelectric dipole moment (qCEDM) operators $\left[C_{uG}\right]_{11}$, $\left[C_{dG}\right]_{11,22}$ are at the moment only preliminary Bhattacharya:2023qwf ; Abramczyk:2017oxr ; Kim:2021qae . Currently, the best estimates come from QCD sum rules Pospelov:2000bw ; Pospelov:2005pr ; Lebedev:2004va ; Fuyuto:2012yf . In Refs. Pospelov:2000bw ; Pospelov:2005pr ; Lebedev:2004va , the $u$ and $d$ qCEDM contribution to the neutron and proton EDMs was estimated to be
@@
with roughly 50% uncertainty. Here $g_{s}$ is the strong coupling constant. We will use the estimates in Eq. (37) for our numerics. Ref. Fuyuto:2012yf finds a larger value and a substantial contribution from the strange qCEDM
@@
Estimates based on the dominance of long-distance contributions mediated by pion-nucleon couplings end up in a similar range Seng:2014pba ; deVries:2016jox , though, of course, they are affected by uncontrolled systematic errors. The above matrix elements are given in the assumption that a PQ mechanism cancels the $\bar{\theta}$ term, including the shift in the $\bar{\theta}$ term induced by the chromoelectric operators Pospelov:2005pr . The scale dependence of $g_{s}$ and of the WET coefficients should in principle be compensated by the scale dependence of the hadronic matrix elements. In the sum rule calculations, such scale dependence is not explicitly tracked, and thus, at the moment, the choice of $\mu$ introduces some arbitrariness. Since the conversion between Lattice QCD schemes and $\overline{\rm MS}$ is known Bhattacharya:2015rsa ; Mereghetti:2021nkt , this issue will be overcome once Lattice QCD calculations will become available. The contribution of the Weinberg three-gluon operator $C_{\tilde{G}}$ has been estimated in Refs. Demir:2002gg ; Haisch:2019bml . The two evaluations are in good agreement, and here we use the results of Ref. Haisch:2019bml
@@
The situation for four-fermion operators is even less well-developed. The matching between gradient flow and $\overline{\rm MS}$ renormalization at one loop has been derived in Ref. Buhler:2023gsg . Robust calculations of the nucleon EDM currently do not exist. Here we will use the expression derived at next-to-next-to-leading order (N2LO) in two-flavor chiral perturbation theory Seng:2014pba
@@
with $g_{A}=1.27$. $\kappa_{1,0}$ are the nucleon isovector and isoscalar anomalous magnetic moments, $\kappa_{1}=3.7$, $\kappa_{0}=-0.12$. $\bar{g}_{0}$, $\bar{g}_{1}$, $\bar{d}_{n}$, and $\bar{d}_{p}$ are $PT$-odd low-energy couplings in the nucleon EFT, which implicitly depend on WET Wilson coefficients. In the following, we will provide a recipe to estimate $\bar{g}_{0,1}$ in the case of four-fermion operators.
@@ 3.4.1 Pion nucleon couplings
$P$ and $T$-odd pion-nucleon ($\pi N$) couplings play an important role for both nucleon and nuclear EDMs. In the case of the nucleon EDM, $\pi N$ couplings induce logarithmically-enhanced contributions Crewther:1979pi , as can be seen in Eqs. (42) and (43). Since they are accompanied by unknown low-energy constants, $\pi N$ couplings cannot completely determine the leading contributions to the nucleon EDM, and thus expressions as Eq. (42) will be eventually superseded by Lattice QCD calculations. Knowing the strength of $\bar{g}_{0,1}$, however, can provide guidance for the chiral and momentum extrapolation Dragos:2019oxn and for the removal of excited state contamination Bhattacharya:2021lol in Lattice QCD calculations. Depending on the chiral properties of the WET operators, $\pi N$ couplings furthermore provide one of the leading contributions to the $PT$-odd potential deVries:2012ab ; Maekawa:2011vs ; deVries:2020iea , which then feeds into the calculation of Schiff moments and thus atomic EDMs Haxton:1983dq ; Engel:1999np .
We focus here on the most important $\pi N$ couplings and refer for a more detailed discussion to Ref. deVries:2020iea . We define the $PT$-odd $\pi N$ Lagrangian as
$\mathcal{L}_{\pi N}=\bar{N}\left[\frac{\bar{g}_{0}}{F_{\pi}}\bm{\tau}\cdot\bm{\pi}+\frac{\bar{g}_{1}}{F_{\pi}}\pi_{0}+\frac{\bar{g}_{2}}{F_{\pi}}\left(\pi_{0}\tau_{3}-\frac{1}{3}\bm{\tau}\cdot\bm{\pi}\right)\right]N,$ (44)
where $\bm{\tau}$ are Pauli matrices in isospin space. The couplings in Eq. (44) break chiral symmetry and are thus mostly important for chiral-symmetry-breaking WET operators. These usually come in chiral multiplets, with the CP-even components inducing corrections to the baryon spectrum while the CP-odd components contribute to $\pi N$ couplings. At leading order in chiral perturbation theory, chiral symmetry then provides a relation between corrections to the baryon spectrum and $\pi N$ couplings deVries:2012ab , which survives higher order corrections deVries:2015una ; deVries:2016jox . It can be shown that $\bar{g}_{2}$ only arises at higher order in isospin breaking deVries:2012ab . The couplings $\bar{g}_{0}$ and $\bar{g}_{1}$ receive a “direct” contribution, proportional to the shift in the baryon masses induced by dim-6 operators, and “indirect” contributions, arising from “vacuum alignment” Baluni:1978rf , i.e. from imposing the condition that $\pi_{0}$ and $\eta$ mesons do not disappear into the vacuum in the presence of $PT$-odd interactions. We will thus write
@@
For the qCEDM, the $\pi N$ couplings were extracted using QCD sum rules Pospelov:2001ys ,
@@
Again, we are neglected contributions proportional to $\bar{\theta}_{\rm ind}$, which are canceled if a PQ mechanism is active. In chiral perturbation theory, the indirect part of the coupling is given by deVries:2016jox
@@
The ratio between $\bar{g}_{1}/\bar{g}_{0}$ is determined by the ratio of isoscalar and isovector charges, which, for the value of the sigma term we adopt, is about $17$. The assumption of the dominance of the indirect piece leads to larger values compared to the QCD sum rules calculations, especially in the case of $\bar{g}_{1}$. For the qCEDM operators we will just use the sum rules results, while for the four-fermion operators, where only the indirect pieces can be estimated, we will capture these effects by assigning a large uncertainty. Plugging Eqs. (46) and (47) or Eqs. (48) and (49) in the chiral perturbation theory expression for the nucleon EDM we recover estimates that agree in magnitude with Eqs. (37)–(39).
We discussed in some detail various estimates for the qCEDM operators to justify our choices for four-fermion operators, for which complete evaluations of the couplings, either on the Lattice or from QCD sum rules are not available. The structure of the couplings is very similar to qCEDM, with a direct piece that could be extracted from dim-6 corrections to the baryon spectrum and an indirect piece determined by the ratio of vacuum matrix elements Cirigliano:2016yhc ; Dekens:2022gha . The operators constructed in Section 2.1.3 belong to the ${\mathbf{8}\times\mathbf{8}}$, $\mathbf{6}\times\mathbf{\bar{6}}$ and $\mathbf{3}\times\mathbf{\bar{3}}$ representations of the $SU(3)$ flavor group Dekens:2022gha . The indirect components of the $\pi N$ couplings are given by
@@
$C^{(1,0)}_{\mathbf{8}\times\mathbf{8}}$ $=\frac{1}{2}\left({\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1221}+{\rm Im}\left[C^{V1,\,LR}_{dd}\right]_{1221}\right)~{},$ (53)
$C^{(8,0)}_{\mathbf{8}\times\mathbf{8}}$ $={\frac{1}{2}}\left({\rm Im}\left[C^{V8,\,LR}_{uddu}\right]_{1221}+{\rm Im}\left[C^{V8,\,LR}_{dd}\right]_{1221}\right)~{},$ (54)
$C^{(1,1)}_{\mathbf{8}\times\mathbf{8}}$ $={\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1111}+\frac{1}{2}{\rm Im}\left[C^{V1,\,LR}_{uddu}\right]_{1221}-\frac{1}{2}{\rm Im}\left[C^{V1,\,LR}_{dd}\right]_{1221}~{},$ (55)